OpenStudy (anonymous):

You need 1 1/4 cups of sugar to make 20 cookies. To make 14 cookies, you will need how many cups of sugar? Where do I begin and how do I solve this? Please help!

hero (hero):

pre-cal is short for pre-calibration

OpenStudy (precal):

create a proportion sugar/cookies 1.25/20=x/14 cross multiply and solve for x 14(1.25)=20x divide by 20 .875 or 7/8 = x

OpenStudy (precal):

No hero it is for precalculus

hero (hero):

I know. I just wanted to say something off-topic

OpenStudy (precal):

ok lets help them out, especially since the first posted solution is incorrect.

OpenStudy (anonymous):

so the answer 7/8?

OpenStudy (anonymous):

No the answer was 7/8 It was right

OpenStudy (precal):

I believe this is a simple proportion problem. Let's look at if the solution is reasonable. You need 1.25 cups of sugar to make 20 cookies. If 1.92500 is correct, why would you need more sugar to make less cookies. Note: we are trying to make 14 cookies.

OpenStudy (anonymous):

nice try blue

OpenStudy (anonymous):

very funny